OpenStudy (anonymous):

Find the point in which the line that passes through the points (1,0,1) and (4,-2,2) intersects the plane x+y+z=6

OpenStudy (jamesj):

Write the equation for the line in parametric form. Then solve for when the ordinates of that form sum to 6.

OpenStudy (anonymous):

okay thx dude

OpenStudy (slaaibak):

would one of the parametric forms be: x = 1 + 3t y = -2t z = 1 + t ? I also need to study this some time..

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

so we set the parametric equation = to 6 and solve?

OpenStudy (anonymous):

or set t=6?

OpenStudy (jamesj):

You set x + y + z = 6 and solve

OpenStudy (jamesj):

Which makes sense as it is the only other piece of information you actually have in the problem.

OpenStudy (anonymous):

so (5/3, -3, 5)?

OpenStudy (anonymous):

-2 sorry

OpenStudy (anonymous):

(5/3, -2, 5)

OpenStudy (jamesj):

What are your equations for the parametric form of the line?

OpenStudy (anonymous):

i'm using the parametric equations above...then setting 6 = 1+3t, 6= -2t, 6 = 1+t and solving each for t

OpenStudy (jamesj):

No, you're not thinking.

OpenStudy (jamesj):

For each value of t (x,y,z) = (1+3t, -2t, 1+t) is a point on the line.

OpenStudy (jamesj):

it cannot be that t is different for each ordinate; i.e., x doesn't have one value of t and y another etc.

OpenStudy (jamesj):

Do you understand where that parametric form comes from?

OpenStudy (anonymous):

okay...i just figured it out lol

OpenStudy (jamesj):

So what's your final answer?

OpenStudy (anonymous):

hang on i'm writing it out

OpenStudy (anonymous):

on paper first

OpenStudy (anonymous):

okay..so the line intersects the xy plane when z = 6 correct?

OpenStudy (jamesj):

No. I recommend you go back and read the section in your book on the parametric form of lines.

OpenStudy (anonymous):

or z = 0

OpenStudy (jamesj):

No.

OpenStudy (anonymous):

u get anywhere with it ska?

OpenStudy (slaaibak):

1+3t -2t + 1+t =6 ? Then solve for t. when you solve t, plug it into this: (1+3t, -2t, 1+t)

OpenStudy (anonymous):

oh i have no idea...was just asking if you got anywhere with it...a shame he's the only one that can help with calc 3

OpenStudy (jamesj):

What slaaibak suggests is exactly right. You need to convince yourself that this makes sense.